# Metric Distortion in Peer Selection

### v1

@@ 3.2 Utilitarian $q$-Cost
Similarly to the classic setting with disjoint voters and candidates, the distortion of voting rules heavily depends on the value of $q$. Indeed, a result by Caragiannis et al. (2022b) directly implies the existence of $(n,k)$-voting rules with distortion $3$ for $q$-cost whenever $q>\frac{k}{2}$, since their result holds in a more general setting with disjoint voters and candidates and general distance metrics. We complement this result in Section 3.2.1 by providing a lower bound that ranges from $\frac{3}{2}$ and $2$ as $q$ varies between $\big{\lceil}\frac{k}{2}\big{\rceil}+1$ and $k$. For $q\leq\frac{k}{2}$, Caragiannis et al. show that no rule provides a bounded distortion. We show that this impossibility still holds in our setting. In Section 3.2.2, we study the case where $q=k=2$ in further detail and achieve the best-possible distortions of $\frac{4}{3}$ and $2$ for odd and even $n$, respectively, through natural voting rules that are able to leverage the different objectives involved in the problem.
@@ 3.2.1 Impossibility Results
Formally, we first consider the case with $B\subseteq S$ and define the distance metric $d_{1}$ on $A$ given by the following positions $x\in(-\infty,\infty)^{n}$: $x_{a}=i-1$ for every $a\in A_{i}$ and every $i\in[p]$, and $x_{a}=2(p-1)$ for every $a\in B$. It is not hard to see that $d_{1}\ \rhd\succ$; see Figure 2 for an illustration. Since $B\subseteq S$, we have that $\big{|}S\cap\bigcup_{i\in[p]}A_{i}\big{|}\leq k-q$. Hence, from an averaging argument, there exists $j\in[p]$ with
@@
where we used our assumption $n\geq 2k+q$. From the definition of $q$-cost, we have that $\text{SC}(S,a;d_{1})=0$ for every $a\in A_{i}$ and every $i\in[p]$. For each $a\in B$, we have $\text{SC}(S,a;d_{1})=p-1$. Combining these facts with eq. 9, we obtain
@@
We now consider the case with $B\not\subseteq S$ and define the distance metric $d_{2}$ on $A$ given by the following positions $x\in(-\infty,\infty)^{n}$: $x_{a}=i-1$ for every $a\in A_{i}$ and every $i\in[p]$, and $x_{a}=p-1+Mn$ for every $a\in B$. It is not hard to see that $d_{2}\ \rhd\succ$; see Figure 2 for an illustration. Since $B\not\subseteq S$, we have that $|S\cap B|<q$ and thus, by the definition of $q$-cost, we have
@@
On the other hand, consider the set $T=B\cup\cup_{i\in[p-1]}T_{i}$, where $T_{i}\subseteq A_{i}$ and $|T_{i}|\geq q$ for every $i\in[p-1]$. Note that this set exists because $(p-1)q=k-q$ and
@@
where we used our assumption $n\geq 2k+q$. From the definition of $q$-cost, we have that $\text{SC}(T,a;d_{2})=0$ for every $a\in A_{i}$ and every $i\in[p-1]$ and $\text{SC}(T,a;d_{2})=0$ for every $a\in B$. For each $a\in A_{p}$, we have $\text{SC}(T,a;d_{2})=1$. Combining these facts with eq. 10, we obtain
@@
We let $k$ and $q$ be as in the statement and fix $n=2(3q-k-2)$, and consider an arbitrary $(n,k)$-voting rule $f$. We partition the agents into four sets $A=\dot{\bigcup}_{i=1}^{4}A_{i}$ such that $|A_{1}|=|A_{4}|=q-1$ and $|A_{2}|=|A_{3}|=2q-k-1$. Note that all these values lie between $1$ and $q-1$. Indeed, this is trivial for $|A_{1}|$ and $|A_{4}|$, whereas for $|A_{2}|$ and $|A_{3}|$ we have $12q-k-1\geq 2\big{(}\frac{k}{2}+1\big{)}-k-1=$ 1 and $12q-k-1\leq 2q-q-1=q-$ 1, where we have used that $q$ lies between $\frac{k}{2}+1$ and $k$.
@@
Formally, we first consider the case with $|S\cap(A_{1}\cup A_{2})|<q$ and $|S\cap(A_{3}\cup A_{4})|<q$ and define the distance metric $d_{1}$ on $A$ by the following positions $x\in(-\infty,\infty)^{n}$: $x_{a}=0$ for every $a\in A_{1}\cup A_{2}$ and $x_{a}=2$ for every $a\in A_{3}\cup A_{4}$. It is not hard to check that $d_{1}\ \rhd\succ$; see Figure 3.(b) for an illustration. It is clear that $\text{SC}(S,a;d_{1})=2$ for every $a\in A$. If we consider the alternative committee $S^{\prime}=A_{1}\cup A_{2}\in{A\choose k}$, we have $\text{SC}(S^{\prime},a;d_{1})=0$ for every $a\in A_{1}\cup A_{2}$ and $\text{SC}(S^{\prime},a;d_{1})=2$ for every $a\in A_{3}\cup A_{4}$. We obtain
@@
If $|S\cap(A_{3}\cup A_{4})|\geq q$, we define the distance metric $d_{2}$ on $A$ by the following positions $x\in(-\infty,\infty)^{n}$: $x_{a}=0$ for every $a\in A_{1}\cup A_{2}$, $x_{a}=1$ for every $a\in A_{3}$, and $x_{a}=2$ for every $a\in A_{4}$. It is not hard to check that $d_{2}\ \rhd\succ$; see Figure 3.(b) for an illustration. Since $|S\cap(A_{1}\cup A_{2}\cup A_{3})|\leq(k-q)+|A_{3}|=q-1<q$, we have that $\text{SC}(S,a;d_{2})=2$ for every $a\in A_{1}\cup A_{2}$. Furthermore, since both $|A_{3}|<q$ and $|A_{4}|<q$, we have that $\text{SC}(S,a;d_{2})=1$ for every $a\in A_{3}\cup A_{4}$. If we consider an alternative committee $S^{\prime}\subseteq A_{1}\cup A_{2}\in{A\choose k}$, which exists due to $|A_{1}\cup A_{2}|=3q-k-2\geq q$, we have $\text{SC}(S^{\prime},a;d_{2})=0$ for every $a\in A_{1}\cup A_{2}$, $\text{SC}(S^{\prime},a;d_{2})=1$ for every $a\in A_{3}$, and $\text{SC}(S^{\prime},a;d_{2})=2$ for every $a\in A_{4}$. Thus, we obtain

### v2

@@ 3.1 Utilitarian Additive Cost
In terms of lower bounds, we prove that for any $k\geq 3$, there exists $n$ such that any $(n,k)$-voting rule has distortion at least $1.0914$. To do so, for any fixed $k$ we consider $n$ agents partitioned into sets $A_{1},A_{2},A_{3}$, with $|A_{3}|\leq|A_{1}|\leq k$ and $|A_{2}|+|A_{3}|=k$. We consider two metrics depicted in Figure 2, both with the agents in the same set located at the same point and $x_{a}\leq x_{b}\leq x_{c}$ for any $a\in A_{1},b\in A_{2}$, and $c\in A_{3}$. In one metric, $A_{2}$ is equidistant to $A_{1}$ and $A_{3}$; in the other, $A_{2}$ and $A_{3}$ are located at the same point. Intuitively, a rule performs badly in the first instance if it selects many agents from $A_{3}$ and in the second instance otherwise. We formalize this distinction through a threshold, obtain bounds parameterized on this threshold for each of the corresponding cases, and obtain the overall bound by optimally picking the threshold and the number of agents in each set that maximizes the minimum distortion between the two cases.
Theorem 2.
For every $k\in\mathbb{N}$ with $k\geq 3$, there exists $n\in\mathbb{N}$ with $n\geq k$ such that, for every $(n,k)$-voting rule $f$, $dist(f)\geq 1.0914$ for utilitarian additive social cost.
Proof.
Consider an arbitrary $(n,k)$-voting rule $f$. We partition the agents into three sets $A_{1},A_{2},A_{3}$ with $|A_{1}|=n_{1}$, $|A_{2}|=n_{2}$, and $|A_{3}|=k-n_{2}$. The natural values $n_{1}$ and $n_{2}$ will be fixed later, respecting $n_{2}\leq k$ and $k-n_{2}\leq n_{1}\leq k$. The total number of agents is then $n=n_{1}+k$. We consider the profile $\succ\in\mathcal{L}^{n}(n)$, where $S=f(\succ)$, and
  1. (i)
  2. (ii)
    $b\succ_{a}c$ whenever $a\in A_{2},b\in A_{3},c\in A_{1}$;
In what follows, we distinguish whether or not $f$ selects $k^{\prime}$ or more agents from $A_{3}$, where $k^{\prime}\in[k-n_{2}]$ will be fixed later, and construct appropriate distance metrics to show that, in either case, the distortion is at least the one claimed in the statement. Intuitively, if $f$ selects many agents from $S_{3}$ (more than $k^{\prime}$), we will consider $A_{1}$, $A_{2}$, and $A_{3}$ from left to right, each at a distance of $1$ from each other, so that picking all agents from $A_{2}$ and some agents from $A_{1}$ would lead to a lower or equal social cost (because $|A_{1}|\geq|A_{3}|$). If $f$ selects few agents from $A_{3}$ (less than $k^{\prime}$), we will consider $A_{2}$ and $A_{3}$ at the same location and a distance of $1$ from $A_{1}$, so that picking all $k$ agents from $A_{2}\cup A_{3}$ would lead to a lower or equal social cost (because $|A_{1}|\leq|A_{2}\cup A_{3}|$).
Formally, we first consider the case with $|S\cap A_{3}|\geq k^{\prime}$ and define the distance metric $d_{1}$ on $A$ by the following positions $x\in(-\infty,\infty)^{n}$: $x_{a}=0$ for every $a\in A_{1}$, $x_{a}=1$ for every $a\in A_{2}$, and $x_{a}=2$ for every $a\in A_{3}$. Note that $d_{1}\ \rhd\succ$; see Figure 2 for an illustration. Since $|A_{1}|=n_{1}\geq k-n_{2}=|A_{3}|$ by assumption, defining $T\subset A$ such that $|A_{1}\cap T|=k-k^{\prime}-n_{2}$, $|A_{2}\cap T|=n_{2}$, and $|A_{3}\cap T|=k^{\prime}$, we know that
$\text{SC}(S,A;d_{1})\geq\text{SC}(T,A;d_{1})=n_{1}(n_{2}+2k^{\prime})+n_{2}(k-n_{2})+(k-n_{2})(n_{2}+2(k-k^{\prime}-n_{2})).$
If we consider the alternative committee $S^{\prime}$ such that $|A_{1}\cap S^{\prime}|=k-n_{2}$, $|A_{2}\cap S^{\prime}|=n_{2}$, and $|A_{3}\cap S^{\prime}|=0$, we have $\text{SC}(S^{\prime},a;d_{1})=n_{2}$ for every $a\in A_{1}$, $\text{SC}(S^{\prime},a;d_{1})=k-n_{2}$ for every $a\in A_{2}$, and $\text{SC}(S^{\prime},a;d_{1})=n_{2}+2(k-n_{2})$ for every $a\in A_{3}$. We obtain
$\mathrm{dist}(f(\succ),\mathcal{E})\geq\frac{\text{SC}(S,A;d_{1})}{\text{SC}(S^{\prime},A;d_{1})}$ $\geq\frac{n_{1}(n_{2}+2k^{\prime})+n_{2}(k-n_{2})+(k-n_{2})(n_{2}+2(k-k^{\prime}-n_{2}))}{n_{1}n_{2}+n_{2}(k-n_{2})+(k-n_{2})(n_{2}+2(k-n_{2}))}$
$=1+\frac{2k^{\prime}(n_{1}+n_{2}-k)}{n_{1}n_{2}+n_{2}(k-n_{2})+(k-n_{2})(n_{2}+2(k-n_{2}))}$
$=1+\frac{2k^{\prime}(n_{1}+n_{2}-k)}{n_{1}n_{2}-2kn_{2}+2k^{2}}.$ (9)
Metric $d_{1}$$A_{1}$$A_{2}$$A_{3}$1111Metric $d_{2}$$A_{1}$$A_{2}\cup A_{3}$11
We now consider the case with $|S\cap A_{3}|\leq k^{\prime}-1$ and define the distance metric $d_{2}$ on $A$ by the following positions $x\in(-\infty,\infty)^{n}$: $x_{a}=0$ for every $a\in A_{1}\cup A_{2}$ and $x_{a}=1$ for every $a\in A_{3}$. Note that $d_{2}\ \rhd\succ$; see Figure 2 for an illustration. Since $|A_{1}|=n_{1}\leq k=|A_{2}\cup A_{3}|$ by assumption, defining $T\subset A$ such that $|A_{1}\cap T|=k-(k^{\prime}-1)-n_{2}$ and $|(A_{2}\cup A_{3})\cap T|=n_{2}+(k^{\prime}-1)$, we know that
$\text{SC}(S,A;d_{2})\geq\text{SC}(T,A;d_{2})=n_{1}(n_{2}+(k^{\prime}-1))+k(k-(k^{\prime}-1)-n_{2}).$
If we consider the alternative committee $S^{\prime}$ such that $|A_{1}\cap S^{\prime}|=0$ and $|(A_{2}\cup A_{3})\cap S^{\prime}|=k$, we have $\text{SC}(S^{\prime},a;d_{2})=k$ for every $a\in A_{1}$ and $\text{SC}(S^{\prime},a;d_{2})=0$ for every $a\in A_{2}\cup A_{3}$. We obtain
$\mathrm{dist}(f(\succ),\mathcal{E})\geq\frac{\text{SC}(S,A;d_{2})}{\text{SC}(S^{\prime},A;d_{2})}$ $\geq\frac{n_{1}(n_{2}+(k^{\prime}-1))+k(k-(k^{\prime}-1)-n_{2})}{n_{1}k}$
$=1+\frac{(k-n_{1})(k-n_{2}-k^{\prime}+1)}{kn_{1}}.$ (10)
We conclude that $\text{dist}(f(\succ),\mathcal{E})\geq 1+\gamma$, where $\gamma$ is the optimal value of the following optimization program:
$\max$ $\min\bigg\{\frac{2k^{\prime}(n_{1}+n_{2}-k)}{n_{1}n_{2}-2kn_{2}+2k^{2}},\frac{(k-n_{1})(k-n_{2}-k^{\prime}+1)}{kn_{1}}\bigg\}$ (11)
s.t. $n_{1},n_{2}\leq k,$
$k^{\prime}+n_{2}\leq k,$
$n_{1}+n_{2}\geq k,$
$k^{\prime},n_{1},n_{2}\in\mathbb{N}.$
Intuitively, for large enough $k$ it is not hard to see that the above program has a strictly positive objective value, leading to a distortion strictly greater than one, while for small values of $k$ in can be solved computationally by enumerating all possible solutions. Naturally, the larger the lower bound we choose for $k$, the better the bound is.
Formally, to obtain the claim bound of $1.0914$, it suffices to assume $k\geq 200$. Indeed, for any fixed $k\geq 200$, we take
$n_{1}=\bigg\lceil\frac{3}{5}k\bigg\rceil,\quad n_{2}=\bigg\lceil\frac{13}{20}k\bigg\rceil,\quad k^{\prime}=\bigg\lceil\frac{1}{5}k\bigg\rceil.$
These values trivially satisfy the constraints of the program (11). Moreover, since $k\geq 200$ we have
$n_{1}\leq\frac{3}{5}k+1\leq\frac{121}{200}k,\quad n_{2}\leq\frac{13}{20}k+1\leq\frac{131}{200}k,\quad k^{\prime}\leq\frac{1}{5}k+1\leq\frac{41}{200}k.$ (12)
Therefore,
$\frac{2k^{\prime}(n_{1}+n_{2}-k)}{n_{1}n_{2}-2kn_{2}+2k^{2}}\geq\frac{2\cdot\frac{1}{5}k\big(\frac{3}{5}k+\frac{13}{20}k-k)}{\frac{121}{200}k\cdot\frac{13}{20}k-2k\cdot\frac{13}{20}k+2k^{2}}=\frac{\frac{2}{5}\cdot\frac{1}{4}}{\frac{1573-5200+8000}{4000}}=\frac{400}{4373}\approx 0.09147,$
where the first inequality follows from (12) and the fact that $n_{1}n_{2}-2kn_{2}$ is decreasing in $n_{2}$ since $n_{1}<2k$. On the other hand,
$\frac{(k-n_{1})(k-n_{2}-k^{\prime}+1)}{kn_{1}}\geq\frac{(k-\frac{121}{200}k)(k-\frac{131}{200}k-\frac{41}{200}k)}{\frac{121}{200}k^{2}}=\frac{\frac{79}{200}\cdot\frac{28}{200}}{\frac{121}{200}}=\frac{2212}{24200}\approx 0.09140,$
where the first inequality again follows from (12). Combining these two inequalities, we obtain $\text{dist}(f(\succ),\mathcal{E})>1.0914$, concluding the proof for $k\geq 200$.
For values $k\leq 200$, we include in Table 2 feasible (and even optimal) solutions for the program (11) alongside their objective value. Since all objective values are strictly greater than $0.0914$, this completes the proof. ∎
$k$ $(n_{1},n_{2},k^{\prime})$ $\gamma$ $k$ $(n_{1},n_{2},k^{\prime})$ $\gamma$ $k$ $(n_{1},n_{2},k^{\prime})$ $\gamma$ $k$ $(n_{1},n_{2},k^{\prime})$ $\gamma$ 33 $(2,2,1)$ $0.1667$ $53$ $(30,37,10)$ $0.0998$ $103$ $(63,67,21)$ $0.0974$ $153$ $(93,99,32)$ $0.0970$ 44 $(2,3,1)$ $0.1429$ $54$ $(32,36,11)$ $0.0995$ $104$ $(62,68,22)$ $0.0977$ $154$ $(90,104,30)$ $0.0969$ 55 $(3,3,2)$ $0.1333$ $55$ $(34,36,11)$ $0.0996$ $105$ $(64,68,22)$ $0.0976$ $155$ $(92,104,30)$ $0.0969$ 66 $(3,5,1)$ $0.1481$ $56$ $(33,37,12)$ $0.0996$ $106$ $(61,72,21)$ $0.0974$ $156$ $(94,103,31)$ $0.0970$ 77 $(3,6,1)$ $0.1250$ $57$ $(35,36,13)$ $0.0992$ $107$ $(63,72,21)$ $0.0978$ $157$ $(96,103,31)$ $0.0970$ 88 $(4,6,2)$ $0.1250$ $58$ $(35,38,12)$ $0.0986$ $108$ $(65,72,21)$ $0.0978$ $158$ $(98,101,33)$ $0.0969$ 99 $(4,7,2)$ $0.1250$ $59$ $(34,40,12)$ $0.0997$ $109$ $(67,71,22)$ $0.0978$ $159$ $(95,106,31)$ $0.0967$ $10$ $(6,7,2)$ $0.1176$ $60$ $(36,40,12)$ $0.1000$ $110$ $(64,75,21)$ $0.0974$ $160$ $(97,105,32)$ $0.0967$ $11$ $(5,8,3)$ $0.1091$ $61$ $(38,39,13)$ $0.0992$ $111$ $(66,74,22)$ $0.0974$ $161$ $(99,105,32)$ $0.0968$ $12$ $(7,9,2)$ $0.1185$ $62$ $(35,43,12)$ $0.0995$ $112$ $(68,74,22)$ $0.0975$ $162$ $(101,104,33)$ $0.0969$ $13$ $(5,11,2)$ $0.1121$ $63$ $(37,43,12)$ $0.0992$ $113$ $(70,73,23)$ $0.0975$ $163$ $(100,104,35)$ $0.0966$ $14$ $(9,9,3)$ $0.1086$ $64$ $(39,42,13)$ $0.0992$ $114$ $(69,73,25)$ $0.0973$ $164$ $(97,109,33)$ $0.0969$ $15$ $(8,11,3)$ $0.1154$ $65$ $(41,41,14)$ $0.0991$ $115$ $(69,76,23)$ $0.0971$ $165$ $(99,110,32)$ $0.0970$ $16$ $(10,11,3)$ $0.1111$ $66$ $(38,45,13)$ $0.0986$ $116$ $(68,77,24)$ $0.0974$ $166$ $(101,108,34)$ $0.0969$ $17$ $(9,12,4)$ $0.1046$ $67$ $(40,45,13)$ $0.0986$ $117$ $(70,77,24)$ $0.0976$ $167$ $(103,107,35)$ $0.0967$ $18$ $(9,14,3)$ $0.1111$ $68$ $(42,44,14)$ $0.0986$ $118$ $(72,76,25)$ $0.0975$ $168$ $(100,112,33)$ $0.0967$ 1919 $(11,13,4)$ $0.1078$ $69$ $(41,46,14)$ $0.0990$ $119$ $(69,81,23)$ $0.0974$ $169$ $(102,112,33)$ $0.0968$ 2020 $(13,13,4)$ $0.1069$ $70$ $(43,45,15)$ $0.0987$ $120$ $(71,81,23)$ $0.0974$ $170$ $(104,111,34)$ $0.0968$ 2121 $(12,14,5)$ $0.1071$ $71$ $(40,48,15)$ $0.0982$ $121$ $(73,80,24)$ $0.0974$ $171$ $(106,110,35)$ $0.0968$ $22$ $(12,16,4)$ $0.1053$ $72$ $(42,49,14)$ $0.0991$ $122$ $(75,80,24)$ $0.0975$ $172$ $(103,114,34)$ $0.0965$ $23$ $(14,15,5)$ $0.1038$ $73$ $(44,49,14)$ $0.0989$ $123$ $(77,78,26)$ $0.0971$ $173$ $(105,114,34)$ $0.0966$ $24$ $(13,17,5)$ $0.1058$ $74$ $(46,47,16)$ $0.0987$ $124$ $(74,83,24)$ $0.0971$ $174$ $(104,114,36)$ $0.0967$ $25$ $(15,17,5)$ $0.1067$ $75$ $(43,52,14)$ $0.0985$ $125$ $(76,82,25)$ $0.0972$ $175$ $(109,113,35)$ $0.0967$ $26$ $(15,18,5)$ $0.1020$ $76$ $(45,51,15)$ $0.0984$ $126$ $(78,82,25)$ $0.0972$ $176$ $(103,118,35)$ $0.0966$ $27$ $(14,20,5)$ $0.1032$ $77$ $(47,51,15)$ $0.0984$ $127$ $(77,82,27)$ $0.0971$ $177$ $(105,119,34)$ $0.0968$ $28$ $(16,20,5)$ $0.1042$ $78$ $(49,50,16)$ $0.0986$ $128$ $(74,86,26)$ $0.0969$ $178$ $(107,118,35)$ $0.0968$ $29$ $(18,19,6)$ $0.1041$ $79$ $(48,51,17)$ $0.0981$ $129$ $(76,86,26)$ $0.0973$ $179$ $(109,118,35)$ $0.0968$ $30$ $(17,20,7)$ $0.1020$ $80$ $(48,53,16)$ $0.0979$ $130$ $(78,86,26)$ $0.0974$ $180$ $(111,116,37)$ $0.0967$ $31$ $(17,22,6)$ $0.1030$ $81$ $(47,54,17)$ $0.0982$ $131$ $(80,85,27)$ $0.0973$ $181$ $(108,121,35)$ $0.0966$ $32$ $(19,22,6)$ $0.1021$ $82$ $(49,54,17)$ $0.0986$ $132$ $(77,90,25)$ $0.0971$ $182$ $(110,120,36)$ $0.0966$ $33$ $(21,21,7)$ $0.1022$ $83$ $(51,53,18)$ $0.0983$ $133$ $(79,89,26)$ $0.0971$ $183$ $(112,120,36)$ $0.0967$ $34$ $(20,23,7)$ $0.1029$ $84$ $(48,58,16)$ $0.0982$ $134$ $(81,89,26)$ $0.0972$ $184$ $(114,119,37)$ $0.0967$ $35$ $(22,22,8)$ $0.1013$ $85$ $(50,57,17)$ $0.0983$ $135$ $(83,88,27)$ $0.0972$ $185$ $(116,117,39)$ $0.0965$ $36$ $(22,24,7)$ $0.1006$ $86$ $(52,57,17)$ $0.0983$ $136$ $(85,86,29)$ $0.0971$ $186$ $(110,123,38)$ $0.0966$ $37$ $(21,26,7)$ $0.1029$ $87$ $(54,56,18)$ $0.0983$ $137$ $(82,91,27)$ $0.0969$ $187$ $(112,123,38)$ $0.0967$ $38$ $(23,25,8)$ $0.1024$ $88$ $(51,60,17)$ $0.0979$ $138$ $(84,91,27)$ $0.0969$ $188$ $(114,122,39)$ $0.0967$ $39$ $(25,24,9)$ $0.1005$ $89$ $(53,60,17)$ $0.0978$ $139$ $(83,91,29)$ $0.0971$ $189$ $(111,127,37)$ $0.0966$ $40$ $(22,28,8)$ $0.1015$ $90$ $(55,59,18)$ $0.0979$ $140$ $(88,90,28)$ $0.0971$ $190$ $(113,127,37)$ $0.0966$ $41$ $(24,28,8)$ $0.1013$ $91$ $(57,59,18)$ $0.0980$ $141$ $(82,95,28)$ $0.0970$ $191$ $(115,126,38)$ $0.0966$ $42$ $(26,27,9)$ $0.1009$ $92$ $(56,59,20)$ $0.0978$ $142$ $(84,95,28)$ $0.0971$ $192$ $(117,126,38)$ $0.0967$ $43$ $(23,31,8)$ $0.1009$ $93$ $(56,62,18)$ $0.0974$ $143$ $(86,95,28)$ $0.0972$ $193$ $(119,125,39)$ $0.0967$ $44$ $(27,28,10)$ $0.1002$ $94$ $(55,63,19)$ $0.0981$ $144$ $(88,94,29)$ $0.0972$ $194$ $(116,129,38)$ $0.0965$ $45$ $(27,30,9)$ $0.1000$ $95$ $(57,63,19)$ $0.0982$ $145$ $(90,92,31)$ $0.0969$ $195$ $(118,129,38)$ $0.0965$ $46$ $(26,31,10)$ $0.1003$ $96$ $(59,62,20)$ $0.0980$ $146$ $(87,97,29)$ $0.0969$ $196$ $(120,128,39)$ $0.0965$ $47$ $(28,31,10)$ $0.1011$ $97$ $(56,66,19)$ $0.0978$ $147$ $(89,97,29)$ $0.0969$ $197$ $(122,128,39)$ $0.0966$ $48$ $(30,30,11)$ $0.1000$ $98$ $(58,66,19)$ $0.0978$ $148$ $(91,96,30)$ $0.0970$ $198$ $(124,125,42)$ $0.0964$ $49$ $(27,34,10)$ $0.0998$ $99$ $(60,65,20)$ $0.0978$ $149$ $(93,95,31)$ $0.0970$ $199$ $(118,132,40)$ $0.0966$ $50$ $(29,34,10)$ $0.1005$ $100$ $(62,65,20)$ $0.0979$ $150$ $(92,95,33)$ $0.0967$ $200$ $(120,133,39)$ $0.0967$ $51$ $(31,34,10)$ $0.1004$ $101$ $(59,68,20)$ $0.0974$ $151$ $(89,100,31)$ $0.0969$ $52$ $(33,33,11)$ $0.0997$ $102$ $(61,68,20)$ $0.0974$ $152$ $(91,100,31)$ $0.0970$
Table 2: Optimal solutions for the program (11) for all integer values of $k$ between $3$ and $200$, alongside their objective value.
@@ 3.2 Utilitarian $q$-Cost
Similarly to the setting with disjoint voters and candidates, the distortion of voting rules heavily depends on the value of $q$. The key intuition is that when $q$ is large, the $q$-cost objective emphasizes the lower-ranked positions in each agent’s ranking, which leads to a higher optimal cost. This makes the problem easier, as the selected committee can better approximate the worst-case distances. In contrast, when $q$ is small, the objective focuses on the top few distances, so there may exist very good committees that are difficult to identify with only ordinal information, leading to greater potential distortion. Indeed, a result by Caragiannis et al. (2022b) implies the existence of $(n,k)$-voting rules with distortion $3$ for $q$-cost whenever $q>\frac{k}{2}$, since their result holds in the setting with disjoint voters and candidates and general distance metrics. We complement this result with a lower bound that ranges from $\frac{3}{2}$ to $2$ as $q$ varies between $\big\lceil\frac{k}{2}\big\rceil+1$ and $k$. For $q\leq\frac{k}{2}$, Caragiannis et al. (2022b) showed that no rule provides a bounded distortion; we prove that this still holds in our setting.
In Section 3.2.2, we study the case where $q=k=2$ in further detail and achieve the best-possible distortions of $\frac{4}{3}$ and 22 for odd and even $n$, respectively, through natural voting rules that are able to leverage the different objectives involved in the problem.
@@ 3.2.1 Impossibility Results
Formally, we first consider the case with $B\subseteq S$ and define the distance metric $d_{1}$ on $A$ given by the following positions $x\in(-\infty,\infty)^{n}$: $x_{a}=i-1$ for every $a\in A_{i}$ and every $i\in[p]$, and $x_{a}=2(p-1)$ for every $a\in B$. Note that $d_{1}\ \rhd\succ$; see Figure 3 for an illustration. Since $B\subseteq S$, we have that $\big|S\cap\bigcup_{i\in[p]}A_{i}\big|\leq k-q$. Hence, from an averaging argument, there exists $j\in[p]$ with
@@
where we used our assumption $n\geq 2k+q$. From the definition of $q$-cost, we have that $\text{SC}(S,a;d_{1})=0$ for every $a\in A_{i}$ and every $i\in[p]$. For each $a\in B$, we have $\text{SC}(S,a;d_{1})=p-1$. Combining these facts with inequality 13, we obtain
@@
We now consider the case with $B\not\subseteq S$ and define the distance metric $d_{2}$ on $A$ given by the following positions $x\in(-\infty,\infty)^{n}$: $x_{a}=i-1$ for every $a\in A_{i}$ and every $i\in[p]$, and $x_{a}=p-1+Mn$ for every $a\in B$. Note that $d_{2}\ \rhd\succ$; see Figure 3 for an illustration. Since $B\not\subseteq S$, we have that $|S\cap B|<q$ and thus, by the definition of $q$-cost, we have
@@
On the other hand, consider the set $T=B\cup\cup_{i\in[p-1]}T_{i}$, where $T_{i}\subseteq A_{i}$ and $|T_{i}|\geq q$ for every $i\in[p-1]$. Note that this set exists because (p1)q=kq(p-1)q=k-q and
@@
where we used our assumption $n\geq 2k+q$. From the definition of $q$-cost, we have that $\text{SC}(T,a;d_{2})=0$ for every $a\in A_{i}$ and every $i\in[p-1]$ and $\text{SC}(T,a;d_{2})=0$ for every $a\in B$. For each $a\in A_{p}$, we have $\text{SC}(T,a;d_{2})=1$. Combining these facts with inequality 14, we obtain
@@
We let $k$ and $q$ be as in the statement and fix n=2(3qk2)n=2(3q-k-2), and consider an arbitrary $(n,k)$-voting rule $f$. We partition the agents into four sets $A=\dot{\bigcup}_{i=1}^{4}A_{i}$ such that $|A_{1}|=|A_{4}|=q-1$ and $|A_{2}|=|A_{3}|=2q-k-1$. Note that all these values lie between $1$ and q1q-1. Indeed, this is trivial for $|A_{1}|$ and $|A_{4}|$, whereas for $|A_{2}|$ and $|A_{3}|$ we have $2q-k-1\geq 2\big(\frac{k}{2}+1\big)-k-1=1$ and $2q-k-1\leq 2q-q-1=q-1$, where we have used that $q$ lies between $\frac{k}{2}+1$ and $k$.
@@
Formally, we first consider the case with $|S\cap(A_{1}\cup A_{2})|<q$ and $|S\cap(A_{3}\cup A_{4})|<q$ and define the distance metric $d_{1}$ on $A$ by the following positions $x\in(-\infty,\infty)^{n}$: $x_{a}=0$ for every $a\in A_{1}\cup A_{2}$ and $x_{a}=2$ for every $a\in A_{3}\cup A_{4}$. Note that $d_{1}\ \rhd\succ$; see Figure 4.(b) for an illustration. It is clear that $\text{SC}(S,a;d_{1})=2$ for every $a\in A$. If we consider the alternative committee $S^{\prime}=A_{1}\cup A_{2}\in{A\choose k}$, we have $\text{SC}(S^{\prime},a;d_{1})=0$ for every $a\in A_{1}\cup A_{2}$ and $\text{SC}(S^{\prime},a;d_{1})=2$ for every $a\in A_{3}\cup A_{4}$. We obtain
@@
If $|S\cap(A_{3}\cup A_{4})|\geq q$, we define the distance metric $d_{2}$ on $A$ by the following positions $x\in(-\infty,\infty)^{n}$: $x_{a}=0$ for every $a\in A_{1}\cup A_{2}$, $x_{a}=1$ for every $a\in A_{3}$, and $x_{a}=2$ for every $a\in A_{4}$. Note that $d_{2}\ \rhd\succ$; see Figure 4.(b) for an illustration. Since $|S\cap(A_{1}\cup A_{2}\cup A_{3})|\leq(k-q)+|A_{3}|=q-1<q$, we have that $\text{SC}(S,a;d_{2})=2$ for every $a\in A_{1}\cup A_{2}$. Furthermore, since both $|A_{3}|<q$ and $|A_{4}|<q$, we have that $\text{SC}(S,a;d_{2})=1$ for every $a\in A_{3}\cup A_{4}$. If we consider an alternative committee $S^{\prime}\subseteq A_{1}\cup A_{2}\in{A\choose k}$, which exists due to $|A_{1}\cup A_{2}|=3q-k-2\geq q$, we have $\text{SC}(S^{\prime},a;d_{2})=0$ for every $a\in A_{1}\cup A_{2}$, $\text{SC}(S^{\prime},a;d_{2})=1$ for every $a\in A_{3}$, and $\text{SC}(S^{\prime},a;d_{2})=2$ for every $a\in A_{4}$. Thus, we obtain